OpenStudy (anonymous):

Given the following functions f(x) and g(x), solve for (f ⋅ g)(2) and select the correct answer below: f(x) = 3x2 + 2 g(x) = x − 8 −42 −84 42 84

OpenStudy (taramgrant0543664):

So your first step is to multiply the function f by function g (3x2+2)(x-8)=?

OpenStudy (taramgrant0543664):

You have to expand first (3x^2+2)(x-8)=? = 3x^3-24x^2+2x-16

OpenStudy (taramgrant0543664):

Then you can substitute in 2 as x

OpenStudy (taramgrant0543664):

3(2)^3-24(2)^2-2(2)-16=? What did you get?

OpenStudy (taramgrant0543664):

I got -84 You can tell this also because 3x+2 is 14 and x-8 is -6 which equals -84